OpenStudy (anonymous):

Please Help! simplify √-100 these are my options -10 -10i 10i 10

OpenStudy (anonymous):

10i

OpenStudy (anonymous):

\[\sqrt{-100}=\sqrt{100}\sqrt{-1}=10i\]

OpenStudy (anonymous):

Thank you so much !